the two numbers sum is 1000 and their difference is 998.the two numbers are?
Answer by Cromlix(4381)

x + y = 1000
x - y = 998
Add
2x = 1998
x = 999
Substitute x = 999
into:
x + y = 1000
999 + y = 1000
y = 1000 - 999
y = 1
Two numbers: 1 and 999